OpenStudy (anonymous):

A box contains 20 slips of paper. 8 of them are blue and are number from 1 to 8. 12 of them are pink and are number from 9 to 20. a) Find the probability of selecting a pink card or a multiple of 4. b) Find the probability of selecting a blue card or a card with the number 18 on it.

OpenStudy (anonymous):

Multiples of 4: 4, 8, 12, 16, 20

OpenStudy (anonymous):

Probability of selecting a pink card: \[\frac{ 12 }{ 20 }\]Probability of selecting a multiple of 4: \[\frac{ 5 }{ 20 }\]Multiply the 2 probabilities to get the one probability:\[\frac{ 12 }{ 20 }\times \frac{ 5 }{ 20 }=\frac{ 3 }{ 20 }\]

OpenStudy (anonymous):

Probability of selecting a blue card: \[\frac{ 8 }{ 20 }\]Probability of selecting the card with 18 on it: \[\frac{ 1 }{ 20 }\]Multiply the 2 probabilities to get one probability: \[\frac{ 8 }{ 20 }\times \frac{ 1 }{ 20 }=\frac{ 1 }{ 50 }\]

OpenStudy (anonymous):

Thank you so much & for explaining it! :)

OpenStudy (anonymous):

No problem. :)